# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ 4.3 Event-Study Analysis: Response to the Advers Health Shocks
Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3. If the negative impacts of the health shocks on consumption were derived from the income losses, the results from this structural estimation approach are similar to those in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. The results are materially similar to the reduced-form results in Column (1), and the weak instrument issues are not relevant in most of the regressions. Overall, the estimates are systematically larger than those reported in Table 4. This partially reflects that the endogeneity is mitigated using the instrumental variable approach. It is important to remember that, however, this system intends to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks. In other words, I do not explicitly take the exogenous favorable shocks into account in the reduced-form equation. This would be another explanation for the difference in the results from my baseline results (Table 4). For example, expenses on the allowance for children could be easily cut during economic hardship but may not respond to favorable shocks. The expenses for furniture are also likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events in my dataset. [^67] In other words, the furniture may have been purchased only when the households had extra income. This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^68] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture. While the first-stage $F$-statitics is $38.6$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applicable to the transportation category because the family travels did not happen frequently.
Results: How did the households compensate the losses from the ill?
@@
Net savings 0.385 [0.322] 263 -0.157 [0.116] 263
Net insurance -0.156 [0.197] 263 0.064 [0.091] 263
Net borrowing 0.125 [0.254] 263 -0.051 [0.098] 263
Net credit purchase 1.278 [0.685]* 263 -0.522 [0.214]** 263
Net gifts 0.176 [0.173] 263 -0.072 [0.083] 263
@@
Other member’s earnings 0.003 [0.080] 263 -0.001 [0.033] 263
Sales of assets 0.001 [0.002] 263 -0.001 [0.001] 263
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 2, in which the head’s earning ($\tilde{y}_{i,t}$) is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the “Coef.” columns. Column 2 shows the results of equation 2, in which the head’s earning ($\tilde{y}_{i,t}$) is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $17.5$. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit redeemed minus the credit expenses in the credit purchase. “Net gifts” refer to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
I test the risk-coping strategies for mitigating adverse health shocks in Table 8. Column 1 lists the reduced-form results for the risk-coping strategies in the same manner in Table 5. The estimated coefficients on the net income variables are positive in most cases, suggesting the households’ compensating responses against the shocks. The estimate for the credit purchases category is weakly statistically significant, with the largest magnitude. It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.28$ yen, which accounts for roughly half of their average daily wages ($1.28/2.56$). Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable, which is instrumented by the head’s illness. A similar but much clearer result is obtained: the estimate for the credit purchases category is $-0.52$, which is nearly four times greater magnitude than that reported in Column (1) of Table 5. The estimate indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.52$ yen. This magnitude is consistent with the direct effect of the health shocks on the net credit purchases estimated in Column (1) of Table 5.
@@
Overall, the event-study analysis using adverse health shocks confirms baseline findings and provides a sharper illustration of consumption smoothing behaviors in the case of adverse idiosyncratic shock events. In short, my results show that the unique available response among the THBS households to the idiosyncratic income shock was to use credit purchases. The available compensation via the credit purchases accounts for roughly half of their temporary income losses, which partially contributes to smoothing their food consumption in the short-run, say in a few weeks.
@@
These findings shed new light on the roles of credit institutions among skilled workers’ households during historical industrialization. While the credit purchases could not fully compensate for the losses of their earnings, this implies the importance of the relationships between the workers and the local retail shops. To the best of my knowledge, however, there is little historical work on credit purchases in prewar Japan, and thus, the details of this credit institution are not yet clear. For example, how retailers get the credit information of their customers, whether the contracts are explicit or implicit, and whether the lenders enforce penalties for the debtors. The information on these institutions may be considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, are less likely to be documented officially. Investigating the institutional aspects of credit purchases is another avenue for future research.
@@ 5 Conclusion
This study investigates the consumption behaviors among working-class households in Tokyo. I digitized the daily longitudinal budget survey on the factory worker households conducted in Tsukishima from 1918 to 1920 and implemented a systematic econometric method to obtain comparable estimates for income elasticities. The estimated income elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic adverse income shocks. For consumption subcategories, I found evidence that while expenses on indispensable consumption categories are partially smoothed, expenses on dispensable categories are likely to be cut in response to the income losses.
@@
From a comparative perspective to the other developing economies, the estimated elasticity for total expenditure is slightly higher ($0.56$), although only a few available references exist. The estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$. A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels ($0.53$) is greater than this estimate. Although the measurement errors in aggregate consumption and income in the monthly panel dataset may explain the discrepancy in these estimates, if taken literally, this implies that the Tsukishima factory workers were more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind it will be the subject of future work.
@@
My findings on the attenuation bias due to the measurement errors are potentially significant in that they may prompt a reinterpretation of the results obtained in previous studies. Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure. An important point is that household-level studies often use aggregate data. In this case, modifying the aggregation period to align with wage payments can reduce attenuation. If the aggregation period cannot be changed, clarifying payroll timing and discussing potential bias is useful.
@@
The results from mechanism analyses provide suggestive evidence that credit purchases played an important role in mitigating adverse idiosyncratic shocks to smooth food consumption among households. Although the institutional details of credit purchases should be investigated further, this finding offers a new historical view that building good relationships with local retailers may have been an important strategy for workers to deal with idiosyncratic risk. This implies that understanding the retail system would be important in comprehending regional consumption behavior and, in turn, in understanding the improvement in living standards in prewar Japan.

### 25 Feb 2024 v6

@@ 4.3 Event-Study Analysis: Response to the Advers Health Shocks
Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3. If the negative impacts of the health shocks on consumption were derived from the income losses, the result from the structural estimation approach is similar to that in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. The results are materially similar to the reduced-form results in Column (1). The estimates are systematically larger than those reported in Table 4. While this result shall reflect the bias reduction thanks to the instrumental variable estimation, it is also important to remember that this system is designed to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events. [^67] In other words, the furniture may have been purchased only when the households had extra income. This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^68] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applicable to the transportation category because the family travels did not happen frequently. These cases may be another explanation for the difference in the results from my baseline results in Table 4.
While the estimates for clothes and gifts categories now exceed one, the null of equality to the population mean of one is not rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on the rare health shock events, the findings are materially similar to my baseline results: both categories are sensitive to the shocks as those are dispensable categories.
Results: How did the households compensate for the losses from the ill?
@@
Net savings 0.397 [0.318] 254 -0.167 [0.117] 254
Net insurance -0.168 [0.204] 254 0.071 [0.100] 254
Net borrowing 0.116 [0.261] 254 -0.049 [0.103] 254
Net credit purchase 1.294 [0.674]* 254 -0.546 [0.203]*** 254
Net gifts 0.181 [0.177] 254 -0.077 [0.090] 254
@@
Other member’s earnings 0.001 [0.081] 254 -0.001 [0.034] 254
Sales of assets 0.002 [0.002] 254 -0.001 [0.001] 254
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the subcolumn named “Coef.”. Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earning are listed in the subcolumn named “Coef.”. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
I test the risk-coping strategies for mitigating adverse health shocks in Table 8. Column 1 lists the reduced-form results for the risk-coping strategies in the same manner in Table 5. The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating responses against the shocks. The estimate for the credit purchases category is weakly statistically significant, with the largest magnitude. It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for roughly half of their average daily wages ($1.29/2.56$). Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable. A similar but much clearer result is obtained: the estimate for the credit purchase category is roughly four times greater magnitude than that reported in Column (1) of Table 5. The estimate indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen. This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.696969This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1). The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-coping strategy for dealing with health shocks. Since it shall take more time (steps) from withdrawing and receiving gifts to use the money at the retailers, it is likely that the households prefer to use credit purchases.
@@
Overall, the event-study analysis using adverse health shocks confirms baseline findings and provides a sharper illustration of consumption smoothing behaviors in the case of adverse idiosyncratic shock events. In short, my results show that the unique available response among the THBS households to the idiosyncratic income shock was to use credit purchases. The available compensation via the credit purchases accounts for roughly half of their temporary income losses, which partially contributes to smoothing their food consumption in the short-run.
5 Credit System in Local Economy
@@
Foregoing results shed new light on the roles of credit institutions among skilled workers’ households during historical industrialization. Since the credit purchases are provided by local retailers, this also suggests the importance of the relationships between the workers and retail shops. To the best of my knowledge, however, there is little historical work on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear: for example, how retailers get the credit information of their customers, whether the contracts are explicit or implicit, and whether the lenders enforce penalties for the debtors.
Table 9: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
Selling Item Number of Share of sales Social Class of Sales Method (%) Number of
stores for regular the Regular Customer (%) Cash Credit Monthly competitors
reported customer (%) Upper Middle Other purchase installment reported
White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
Notes: This table summarizes the statistics on the sales in the $326$ retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) are $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel (firewood and charcoal) categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual total sales in 1935. The definition of the social classes (“Upper”; “Middle”; “Other”) is not exactly defined in the report. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey. Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
The information on the credit purchase is considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994), and can hardly be documented. Therefore, systematic information on credit purchases in the prewar period is unavailable. Fortunately, however, a survey report on retail shops in 1935 published by the Tokyo City Chamber of Commerce and Industry remained. Table 9 summarizes the information on the sales in those retail shops of rice, fish, fruit and vegetables, and fuels. [^70] This survey covers 939 shops and was conducted on 31st December 1935. I digitized the statistics for all the shops selling the representative items (rice, fish, fruit and vegetables, and fuels) in the credit purchases. The survey instrument includes detailed questions regarding sales and purchases, making it the most comprehensive of the available documents (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by the scale of businesses. I confirm that the main findings are materially similar. First, many annual sales are from their regular customers, most of whom are from the non-upper classes. For example, $73$% of annual sales in rice retailers are from their regular customers, and more than $80$% of them are from non-upper classes, including the factory workers. Second, and most importantly, credit purchase is the popular sales method, especially among rice and fuel retailers: more than $70$% of the sales were done with credit in both shops. Although cash was also used frequently among greengrocers, $34$% of the sales were settled using credit. This supports the evidence that credit purchase was a popular payment method, and regular customers settled a large part of it. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions in the context of a long-term relationship with their regular customers. The final column of this table further provides suggestive evidence behind this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (roughly $330$ meters) radius. As shown, rice and fuel stores have 2-3 more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
Figure 6: Relationships between Average Credit Purchases and Payments
by Different Time-series Frequencies
(a) Daily panel dataset:
purchase and payment
(b) Daily panel dataset:
net purchase
(c) Semi-monthly panel dataset:
purchase and payment
(d) Semi-monthly panel dataset:
net purchase
Note: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. Figures 6a and 6b show the daily data between 12th January and 11th November 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. All figures are based on the 30 THBS households that used any credit purchases. Source: Created by the author using the THBS sample.
Figure 6: Relationships between Average Credit Purchases and Payments
by Different Time-series Frequencies
Next, I look into consumers’ credit purchase behavior using the THBS dataset. Most of the THBS households ($29/33$) used any credit purchases during their sample period, supporting the evidence that this was a common practice among factory workers in Tsukishima. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the households’ behavior of making the payments for the credit purchase on or soon after their paydays. The purchases series shows a similar tendency but with lower spikes and larger variance. This indicates that the households purchased goods by a certain amount around their paydays and then used it when they needed unconservable food such as fish, fruit, and vegetables. Figure 6b illustrating the net purchases confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figures 6a and 6d show that the average payment exceeds the average purchase in almost all semi-month cells, meaning that many households were able to repay their loans. This is consistent with the findings from the survey report that credit purchases function based on the long-term relationship between retailers and consumers in the local economy.
I then analyze whether the volatility of earnings restricts credit purchases. Since the credit of a worker is unobservable, I use the volatility of earnings as a proxy of credit risk in the following empirical specification:
$I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\mathbf{\chi}+v_{i},$ (4)
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^71] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect is unable to be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. In contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
Table 10: Volatility of Earnings and Use of Credit Purchases
Dependent Variable: Indicator Variable for
Non-CP Users Non-CP + Low-CP Users
(1) (2) (3) (4) (5) (6)
CV of Head’s Earnings $0.003$ $-0.002$ $-0.027$ $0.019$ $0.008$ $-0.005$
$[0.012]$ $[0.014]$ $[0.050]$ $[0.013]$ $[0.017]$ $[0.049]$
Sample mean of the DV $0.12$ $0.12$ $0.12$ $0.24$ $0.24$ $0.24$
Model LPM LPM Probit LPM LPM Probit
Number of THBS households 33 33 33 33 34 33
Family Size Yes Yes Yes Yes Yes Yes
Children aged 6–12 (%) Yes Yes Yes Yes Yes Yes
Children aged 13–16 (%) Yes Yes Yes Yes Yes Yes
Adults aged 17+ (%) Yes Yes Yes Yes Yes Yes
Head’s age No Yes Yes No Yes Yes
Head’s earnings No Yes Yes No Yes Yes
Zero slope ($p$-value) 0.467 0.718 0.336 0.515 0.519 0.451
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets. Notes: The dependent variable in Columns (1)-(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users). The dependent variable in Columns (4)-(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases. Columns (1)-(2) and (4)-(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model. “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings. The sample mean of the CV is $5.76$ (Std. Dev. = $5.72$). All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group). The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6. $F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)-(2) and (4)-(5). $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6). The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al. (1975).
Column (1) of Table 10 presents the result for equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern herein may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households using a relatively smaller amount of credit purchases, say less than the first quartile of the average credit purchases. Columns (4)-(5) summarize the results in the same column layouts as Columns (1)-(3), confirming that the estimates are still close to zero in all the specifications. Finally, the final row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households to use credit purchases. This suggests that retailers may not have put weight on the volatility of earnings when trading with factory workers on credit. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit purchases.
The results so far indicate that retailers have been generous in the earnings volatility risk. This, in turn, implies that retailers may factor risk in the form of interest into their loans or retail prices. Unfortunately, there are no systematic statistics on the interest in credit purchases. Thus, I calculate the risk premiums backward by differencing payments and purchases. This backward calculation technique requires information both on the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were sold in greengrocers (yaoya) in many cases but sometimes sold in variety stores (aramonoya). Fortunately, I found that a specific month in a budget book has documented the price-by-retailer information. [^72] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and variety stores between 1st and 14th February 1919 and paid exactly the same amounts on the 15th. In the latter half of February, the household paid the same amounts to greengrocers and variety stores on 28th, but less amount to the fish stores. Since the budget book for March does not have information on the type of retailers, it is difficult to know exactly how the reimbursement shortfall was handled. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this condition, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. This ease of accessibility to credit purchases might enable workers to smooth consumption in response to adverse idiosyncratic shocks in the short-run.
@@ 6 Conclusion
This study investigates the consumption smoothing behaviors among working-class households in Tokyo. I digitized the daily longitudinal budget survey on the factory worker households conducted in Tsukishima and implemented a systematic econometric method to obtain comparable estimates for income elasticities. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks. Despite this, I found evidence that while expenses on dispensable categories are likely to be cut in response to the income losses, expenses on indispensable categories are partially smoothed.
The estimated elasticity for total expenditure is in a similar range to those in rural economies in developing countries, although only a few available references exist. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$. As for the urban households, the estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, one must be careful when making comparisons with these estimates based on the urban samples, including a variety of occupations. At least, however, the factory workers in the historical metropolis of Japan around WWI had been relatively vulnerable to the idiosyncratic income shocks.
@@
A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. Although the measurement errors in aggregate consumption and income in the monthly panel dataset may explain the discrepancy in these estimates, if taken literally, this implies that the Tsukishima factory workers were more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind the consumptions there would be another subject of future work.
@@
The finding on the attenuation bias due to the measurement errors is potentially important in that they may prompt a reinterpretation of existing results. Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure, albeit the household-level studies often use aggregate data. In this case, modifying the aggregation period to align with wage payments can reduce attenuation bias. Even if the aggregation period is unadjustable, clarifying payroll timing and discussing potential bias is useful.
@@
The results from mechanism analyses provide suggestive evidence that credit purchases played an important role in mitigating adverse idiosyncratic shocks on consumption. Although the institutional details of credit purchases should be investigated further, this finding offers a new historical view of the interdependence between consumers and local retailers. Understanding the local retail system would be important in comprehending regional consumption behavior and, in turn, in understanding the improvement in living standards in prewar Japan.